cindy invested a certain amount of money at 6% interest and 15000 more than that amount at 8% interest her yearly interest was 540 how much did she invest?
Answer by mananth(16949)

cindy invested a certain amount of money at 6% interest and 15000 more than that amount at 8% interest her yearly interest was 5400 how much did she invest?$(x+15000)@8%
$x @ 6%
interest =5400
...
0.06x+0.08(x+15000)=5400
0.06x+0.08x+1200=5400
0.14x=5400-1200
0.14x=4200
x=4200/0.14
x =30,000 @ 6%
$45,000@ 8%
..
m.ananth@hotmail.ca